# BOP Challenge 2023 on Detection, Segmentation and Pose Estimation of Seen and Unseen Rigid Objects

### v1

@@ 4.2 Results on Task 1
Results on the task of 6D object localization of seen objects and properties of the evaluated methods are in Tab. 2. Among the 16 new entries in 2023, three outperform GDRNPP [62, 40], the best method from the 2022 challenge. The best pose estimation pipeline from 2023, GPose2023 [62, 1], is purely learning-based and achieves 85.6 AR, outperforming GDRNPP by 1.9 AR (#1#4 in Tab. 2) with less than half the inference time (2.67 s vs. 6.26 s). GPose2023 deploys the same pose estimation method as GDRNPP but combines it with a more efficient coordinate-guided pose refinement strategy [1] and an improved 2D object detector based on YOLOv8 (see #1#2 in Tab. 4). Without any pose refinement, the RGB-only variants GPose2023-RGB (#21, 72.9 AR) or ZebraPoseSAT-EffnetB4 [54] (#17, 74.9 AR) reach an average inference time of 0.25 seconds per image which are closer to the demands of mobile vision applications. Gains in accuracy are most notable on the industrial ITODD, T-LESS, and HB datasets, whereas on TUD-L and YCB-V we can observe that metrics start to saturate.
@@ 4.3 Results on Tasks 2 and 3
As shown in Tab. 4, GDet2023 [1] based on YOLOv8 [29] achieves 79.8 AP, a moderate +2.5 AP gain over YOLOX [12], the best detector in 2022. YOLOv8 is even less sensitive to the training image domain than YOLOX, achieving 76.9 AP when trained only on synthetic PBR images and neglecting the real training data. In the 2D segmentation of seen objects task (Tab. 5), we see a similar incremental improvement of +3.2 AP achieved by ZebraPoseSAT [54], which predicts object masks from the provided default detections of GDRNPPDet.
@@ 4.4 Results on Task 4
The new task of 6D localization of unseen objects received 14 entries, as presented in Tab. 3. MegaPose [34], a method from 2022, was considered as the baseline and consists of two stages: (1) coarse object pose estimation by finding the rendered template image that is closest to the test image crop, and (2) pose refinement via a render-and-compare strategy. The RBG-only entry Megapose-CNOS_fastSAM+Multih-10 (#9) achieves 54.9 AR and further improves to 62.8 AR by using RGB-D images and an additional refinement with Teaser++ [66], see Megapose-CNOS+Multih_Teaserpp-10 (#3).
GenFlow-MultiHypo16 (#1), the best method for 6D localization of unseen objects, reaches 67.4 AR. This is a remarkable result since the performance is comparable to CosyPose [33], the best method in 6D localization of seen objects from 2020. GenFlow improves the coarse pose estimation stage of MegaPose by running the coarse network in a GMM-based hierarchical manner. For pose refinement, GenFlow adapts the recurrent flow network [14] to also estimate a visibility mask and replaces the pose regression network with a differentiable PnP solver.
Results in Tab. 3 highlight that the runtime is a significant challenge for solving unseen object pose localization. While GenFlow-MultiHypo16 improved the runtime by 4x compared to MegaPose, it still takes 34.58 s per image. SAM6D (#5) [36] based on GeoTransformer [48] is the fastest method by a significant margin with 3.87 s per image while still reaching 61.6 AR (-5.8 AR compared to Genflow-MultiHypo16 #1). Figure 4 shows qualitative comparison of the best method for unseen objects, GenFlow, with the best method for seen objects, GPose.
@@ 4.5 Results on Tasks 5 and 6
2D detection and segmentation of unseen objects in cluttered, occluded environments is a challenging task. Still, as shown in Tab. 6 and Tab. 7, the best method CNOS-FastSAM [44] reaches accuracy of 42.8 mAP in detection and 41.2 mAP in segmentation of unseen objects. For comparison, the instance segmentation accuracy is comparable to Mask R-CNN [17] that reached 40.5 mAP in the BOP challenge 2020 [26] while being trained on more than 1M synthetic and real images of the target objects. CNOS-FastSAM [44] instead relies on DINOv2 [46] features extracted from only 200 rendered reference views per object. All submitted detection and segmentation approaches are RGB-based and rely on SAM-like (Segment Anything) [36] methods to segment object instances in the image.
Despite the substantial progress in unseen object detection and segmentation driven by foundation models, there is still a relatively large gap to methods trained to detect and segment specific objects (compare Tab. 4 and 5). Especially, the amodal detection of occluded instances, i.e., including occluded parts, is a clear challenge for approaches focusing on unseen objects, leading to a gap of 37 mAP between CNOS and GDet2023.
To what extent is this gap in 2D detection performance responsible for the gap in 6D localization of seen and unseen objects? When combined with the default GDRNPPDet detections from Task 2, the best method for 6D localization of unseen objects (GenFlow-MultiHypo16) achieves the pose accuracy of 79.2 AR (#10 Tab. 2). Since this is only -5.9 AR behind GDRNPPDet + GPose2023 (#2), we conclude that better methods for unseen object detection would provide great potential for improving methods for unseen object localization.

### v2

@@ 4.2 Results on Task 1
Results on the task of 6D object localization of seen objects and properties of the evaluated methods are in Tab. 2. Among the 16 new entries in 2023, three outperform GDRNPP [61, 39], the best method from the 2022 challenge. The best pose estimation pipeline from 2023, GPose2023 [61, 67], is purely learning-based and achieves 85.6 ARC, outperforming GDRNPP by 1.9 ARC (#1#4 in Tab. 2) with less than half the inference time (2.67 s vs. 6.26 s). GPose2023 deploys the same pose estimation method as GDRNPP but combines it with a more efficient coordinate-guided pose refinement strategy [67] and an improved 2D object detector based on YOLOv8 (see #1#2 in Tab. 4). Without any pose refinement, the RGB-only variants GPose2023-RGB (#21, 72.9 ARC) or ZebraPoseSAT-EffnetB4 [53] (#17, 74.9 ARC) reach an average inference time of 0.25 seconds per image which are closer to the demands of mobile vision applications. Gains in accuracy are most notable on the industrial ITODD, T-LESS, and HB datasets, whereas on TUD-L and YCB-V we can observe that metrics start to saturate.
@@ 4.3 Results on Tasks 2 and 3
As shown in Tab. 4, GDet2023 [67] based on YOLOv8 [28] achieves 79.8 APC, a moderate +2.5 APC gain over YOLOX [11], the best detector in 2022. YOLOv8 is even less sensitive to the training image domain than YOLOX, achieving 76.9 APC when trained only on synthetic PBR images and neglecting the real training data. In the 2D segmentation of seen objects task (Tab. 5), we see a similar incremental improvement of +3.2 APC achieved by ZebraPoseSAT [53], which predicts object masks from the provided default detections of GDRNPPDet.
@@ 4.4 Results on Task 4
The new task of 6D localization of unseen objects received 14 entries, as presented in Tab. 3. MegaPose [33], a method from 2022, was considered as the baseline and consists of two stages: (1) coarse object pose estimation by finding the rendered template image that is closest to the test image crop, and (2) pose refinement via a render-and-compare strategy. The RGB-only entry Megapose-CNOS_fastSAM+Multih-10 (#9) achieves 54.9 ARC and further improves to 62.8 ARC by using RGB-D images and an additional refinement with Teaser++ [65], see Megapose-CNOS+Multih_Teaserpp-10 (#3).
GenFlow-MultiHypo16 (#1), the best method for 6D localization of unseen objects, reaches 67.4 ARC. This is a remarkable result since the performance is comparable to CosyPose [32], the best method in 6D localization of seen objects from 2020. GenFlow improves the coarse pose estimation stage of MegaPose by running the coarse network in a GMM-based hierarchical manner. For pose refinement, GenFlow adapts the recurrent flow network [13] to also estimate a visibility mask and replaces the pose regression network with a differentiable PnP solver.
Results in Tab. 3 highlight that the run time is a significant challenge for solving unseen object pose localization. While GenFlow-MultiHypo16 improved the run time by 4x compared to MegaPose, it still takes 34.58 s per image. SAM6D (#5) [35] based on GeoTransformer [47] is the fastest method by a significant margin with 3.87 s per image while still reaching 61.6 ARC (-5.8 ARC compared to Genflow-MultiHypo16 #1). Figure 4 shows qualitative comparison of the best method for unseen objects, GenFlow, with the best method for seen objects, GPose.
@@ 4.5 Results on Tasks 5 and 6
2D detection and segmentation of unseen objects in cluttered, occluded environments is a challenging task. Still, as shown in Tab. 6 and Tab. 7, the best method CNOS-FastSAM [43] reaches accuracy of 42.8 mAPC in detection and 41.2 mAP in segmentation of unseen objects. For comparison, the instance segmentation accuracy is comparable to Mask R-CNN [16] that reached 40.5 mAPC in the BOP challenge 2020 [25] while being trained on more than 1M synthetic and real images of the target objects. CNOS-FastSAM [43] instead relies on DINOv2 [45] features extracted from only 200 rendered reference views per object. All submitted detection and segmentation approaches are RGB-based and rely on SAM-like (Segment Anything) [35] methods to segment object instances in the image.
Despite the substantial progress in unseen object detection and segmentation driven by foundation models, there is still a relatively large gap to methods trained to detect and segment specific objects (compare Tab. 4 and 5). Especially, the amodal detection of occluded instances, i.e., including occluded parts, is a clear challenge for approaches focusing on unseen objects, leading to a gap of 37 mAPC between CNOS and GDet2023.
To what extent is this gap in 2D detection performance responsible for the gap in 6D localization of seen and unseen objects? When combined with the default GDRNPPDet detections from Task 2, the best method for 6D localization of unseen objects (GenFlow-MultiHypo16) achieves the pose accuracy of 79.2 ARC (#10 Tab. 2). Since this is only 5.9 ARC behind GDRNPPDet + GPose2023 (#2), we conclude that better methods for unseen object detection would provide great potential for improving methods for unseen object localization.